A woman is 36 and her daughter is 16. At some point during the lives of the two, the woman's will be 3 times the daughter age at that moment. Write an equation which models how old in years the woman will be when she is 3 times as old as the daughter.
Answer by stanbon(75887)

A woman is 36 and her daughter is 16. At some point during the lives of the two, the woman's will be 3 times the daughter age at that moment. Write an equation which models how old in years the woman will be when she is 3 times as old as the daughter.
---------------
Equation:
36+x = 3*(16+x)
36+x = 48+3x
2x = -12
x = -6
-----
Woman's age: 36-6 = 30 yrs
Daughter's age: 16-6= 10 yrs
====================
Cheers,
Stan H.